Determine the standard free energy change for the reaction of hydrazine with hydrogen peroxide to form nitrogen gas and water vapor at 298K.

a. Determine the temperature, in K, at which the reaction of nitrogen gas with oxygen gas to form nitrogen monoxide gas would become spontaneous. ?H°rxn = 180.5kJ, ?S°rxn = 24.9J/K

b. Determine the equilibrium constant for the reaction in part “a” at 298K.

please show me all your work !!!
Answer #1

the temperature, in K, at which the reaction of nitrogen gas with oxygen gas to form nitrogen monoxide gas would become spontaneous

?H°rxn = 180.5kJ, ?S°rxn = 24.9J/K

?G° = ?H° - T?S°

To become spontaneous

?G° = 0

?H° - T?S° = 0

?H° = T?S°

T = ?H° / ?S°

= 180.5*1000 J / (24.9 J/K)

= 7249 K

Part b

?G° = ?H° - T?S°

= 180.5*1000 - 298*24.9

= 173079.8 J

= 173.0798 kJ

?G° = - RT ln K

173079.8 = - 8.314*298*ln K

ln K = - 69.858

Equilibrium constant K = 4.58 x 10^-31